Billiard balls of identical mass roll on a table, and we can ignore the small friction between the balls and the table. The cue ball travels east at a speed of 7.21 m/s and hits the 8 ball which is initially stationary. After the collision the cue ball travels at a speed of 3.05 m/s 29.5 degrees south of its initial eastward heading. What is the direction of the 8 ball in degrees measured counterclockwise from East (+X)? Enter only the number to three significant figures. (Let east be +X, north be +Y, south be -Y, west be -X.)
Answer #1

Given data, m. = mg = m, u = 7.21 m/s, Ug = 0 m/s, v. = 3.05m/s, and 0 = 29.5° Draw a free body diagram for the collision bet= (3.05 m/s)(0.8704)+ Vex Vex = (7.21 m/s)-(2.6546 m/s) = 4.55 m/s Vsy From the law of conservation of momentum in vertical d